OpenStudy (anonymous):

A jar contains 18 apple lollipops, 7 strawberry lollipops, and 10 grape lollipops. Irvin randomly selects a lollipop, gives it to his brother, and then randomly selects another lollipop for himself. What is the probability that Irvin selected a grape lollipop for his brother and an apple lollipop for himself? Write your answer as a percent. Round to the nearest tenth of a percent.

OpenStudy (anonymous):

First choice is "1 of 10 grape lollipops from 35 lollipops" The probablity is 10/35 = 2/7 Second choice is "1 of 18 apple lollipops from 34 lollipops" So probablity is 18/34 = 9/17 Between this is AND so you have to these probablities multiply: P(A) = 2/7 * 9/17 = 18/119 P(A) = 18/119 * 100% = 1800/119 % It is approximaly 15,1 %. Is this right

OpenStudy (anonymous):

probability of getting a grape lollipop:10/35 probability of getting an apple lollipop after grape lollipop has been picked : 18/34 Do 10/35*18/34

OpenStudy (anonymous):

to get 18/119

OpenStudy (anonymous):

which is approximately 15%, so I'd say you are correct

OpenStudy (anonymous):

Thank you